OpenStudy (ranya99):

2. Illinois Jones is being pulled from a snake pit with a rope that breaks if the tension in it exceeds 755N. (a) If Illinois Jones has a mass of 73.0kg and the snake pit is 3.40m deep, what is the minimum time necessary to pull our intrepid explorer from the pit?

OpenStudy (ranya99):

@IrishBoy123 @ospreytriple

OpenStudy (irishboy123):

well, set the force to 755N and go.....!  \(755 - W = ma\) the more he accelerates, the quicker he gets out.

OpenStudy (ranya99):

And he has a mass of 73.0kg

OpenStudy (ranya99):

What about W?

OpenStudy (ranya99):

mg = 73 x 9.8 = 715.4N  break point = 755 N  755 = m x (g + a)  755/m = (g + a)  755/73 = 10.342

OpenStudy (ranya99):

10.342 - g = a = 0.542 m/s^2  3.40/(0.5 x 0.542) = t^2 =12.55 , sq-root = t = 3.5 secs 

OpenStudy (ranya99):

@IrishBoy123 Is my work correct?

OpenStudy (irishboy123):

@Ranya99 hope that worked!!